Let X,...Xn be a random sample from the density fx(x) = 1+θX^θ, 0<x<1

a) Use the Neymar-Pearson lemma to determine the best critical region for testing Ho: θ-θo against H1 θ-θ1 > θo
Answer #1

OLM<] hef f(,0) = (1+0) a wex, o )= (1+0) 1877 ( ) - Lux, Lex,00) dex) = 41400 - 6+0,*T*). - (+9) (2,6,%. BC) (140, 5 (T1 ~